OpenStudy (zootedyoungin):

Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches. The diameter of circle B is 6 inches, the circumference is 18.84 inches, and the area is 28.26 square inches.

OpenStudy (zootedyoungin):

Part A: Using the formula for circumference, solve for the value of pi for each circle. Part B: Use the formula for area and solve for the value of pi for each circle. Part C: What observation can you make about the value of pi for circles A and B?

OpenStudy (welshfella):

Circumference = pi * diameter 21.98 = 7 * pi solve for pi

OpenStudy (zootedyoungin):

solve for pi? pi = 3.14, do i just pretend i dont know what pi is? as if pi were X?

OpenStudy (welshfella):

yea 7 pi = 21.98 divide both sides by 7:- pi = 21.98 / 7

OpenStudy (zootedyoungin):

oh its a one step equation

OpenStudy (welshfella):

right

OpenStudy (zootedyoungin):

lol 21.98/7 = 3.14

OpenStudy (zootedyoungin):

@agent0smith

OpenStudy (zootedyoungin):

@Sachintha

OpenStudy (agent0smith):

What are you stuck on?

OpenStudy (zootedyoungin):

all of it - not pphalk

OpenStudy (zootedyoungin):

i kinda got part a down tho

OpenStudy (agent0smith):

lol i miss pphalke. Part B is almost the same as A, except \[\large Area = \pi r^2 \]

OpenStudy (zootedyoungin):

i find the area of the circles?

OpenStudy (agent0smith):

They gave you the areas "Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches"

OpenStudy (zootedyoungin):

yeah

OpenStudy (agent0smith):

Plug in the area for A (38.465), and radius for A (3.5) into the area formula and solve for pi.

OpenStudy (zootedyoungin):

I got 3.14 for part A and B but dont know how to show the work

OpenStudy (agent0smith):

Circle A 38.465 = pi*3.5^2 Circle B 28.26 = pi*3^2 solve for pi in both

OpenStudy (zootedyoungin):

i dont understand it

OpenStudy (zootedyoungin):

38.465/7

OpenStudy (agent0smith):

38.465 = pi*3.5^2 38.465/(3.5^2) = pi and 28.26 = pi*3^2 28.26/(3^2) = pi

OpenStudy (zootedyoungin):

38.465/7 = 5.495

OpenStudy (agent0smith):

Are you doing part a or part b? The work i showed above is part b

OpenStudy (zootedyoungin):

OHH

OpenStudy (zootedyoungin):

can you help me with part C

OpenStudy (agent0smith):

It's just asking you to comment on the pi you got for part a and b, for the circles

OpenStudy (zootedyoungin):

pi was the same for both circles though

OpenStudy (agent0smith):

Then just say that

OpenStudy (zootedyoungin):

thanks dude